I. A random variable X has pdf: X 0 1 2 f(x) 3 1/8 3/8 3/8 1/8 (a) Find the cumulative distribution function F(x) (b) Graph this distribution function. II. If the probability of a defective bolt is 0.1, find:  (a) the mean, for the number of defective bolts in a total of 400 bolts. (b) the standard deviation, for the number of defective bolts in a total of 400 bolts. 


1
Expert's answer
2021-06-15T14:18:22-0400

I.

a)

"x=0" : cum. freq.=1/8

"x=1" : cum. freq.=1/8+3/8=1/2

"x=2" : cum. freq.=1/8+3/8+3/8=7/8

"x=3" : cum. freq.=1/8+3/8+3/8+1/8=1


b)



II.

a)

"E(X)=0.1\cdot400=40"


b)

"\sigma=\sqrt{npq}=\sqrt{40(1-0.1)}=6"